# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ A.2 Helper Lemmas
If $C_{2}^{\mathsf{C3}}(v)$ is connected to $C_{2}^{\mathsf{C3}}(w)$ in $H_{2}^{\mathsf{C3}}$, then $H_{2}^{\mathsf{C3}}$ is connected. Further, $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is connected and spanning for $G_{1}$. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). We conclude via Lemma 25 that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-3\leq 0$.
Otherwise, that is, $C_{2}^{\mathsf{C3}}(v)$ is not connected to $C_{2}^{\mathsf{C3}}(w)$ in $H_{2}^{\mathsf{C3}}$, Lemma 24 guarantees the existence of an edge $e$ in $G_{2}$ between the vertices of $V(C_{2}^{\mathsf{C3}}(w))$ and the vertices of $V(H_{2}^{\mathsf{C3}}\setminus V(C_{2}^{\mathsf{C3}}(w))$. Therefore, $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Hence, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=3-3=0$.
@@
First, assume that $H_{2}^{\mathsf{C3}}$ is connected. We have that $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is connected and spanning for $G_{1}$. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). We conclude via Lemma 25 that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-4\leq 0$.
@@
If $H_{2}^{\mathsf{C3}}$ is composed of two connected components, assume w.l.o.g. that $C_{2}^{\mathsf{C3}}(u)$ is connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$ and the other connected component in $H_{2}^{\mathsf{C3}}$ is $C_{2}^{\mathsf{C3}}(w)$. In this case, Lemma 24 guarantees that there is an edge $e$ in $E(G_{2})\setminus C_{2}^{\mathsf{C3}}(w)$ between the vertices of $C_{2}^{\mathsf{C3}}(w)$ and the vertices of $V(H_{2}^{\mathsf{C3}})\setminus V(C_{2}^{\mathsf{C3}}(w))$. Therefore, $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Hence, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=3-4\leq 0$.
@@
Otherwise, that is, $C_{2}^{\mathsf{C3}}(u)$, $C_{2}^{\mathsf{C3}}(v)$ and $C_{2}^{\mathsf{C3}}(w)$ are all pairwise disjoint, Lemma 24 guarantees the existence of two edges $F\subseteq E(G_{2})$ such that $H_{2}^{\mathsf{C3}}\cup F$ is connected. Hence, we can analogously to the previous case(s) show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq F\cup\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=4-4=0$. ∎
A.3 Final Lemmas
In this final subsection, we prove Definition 5. The statement is implied by the following three lemmas. We first show that $\mathsf{RED}(G)$ runs in polynomial time. The proof is similar to the corresponding proof in [18].
Lemma 34 (running time).
@@ A.3 Final Lemmas
Let $n=|V(G)|$ and $m=|V(G)|$, and $\sigma=n^{2}+m^{2}$ be the size of the problem. We first argue that every non-recursive step of Algorithms 1 and 2 can be performed in time polynomial in $\sigma$. For all computations performed in Algorithm 1 this follows from [18]. We already argued that the computation in Algorithm 2 in Algorithm 2 can be executed in polynomial time in $\sigma$. Moreover, determining the $\mathsf{C2}$ subcases in Algorithms 2, 2 and 2 and selecting a desired $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution in Algorithms 2 and 2 and $\mathsf{OPT}_{1}^{\mathsf{C3}}$ solution in Algorithm 2 can be done in constant time, as $G_{1}$ is of constant size in this case. Finding the corresponding sets of edges $F$ in Algorithm 2 can also be done in polynomial time in $\sigma$ by enumeration, because Lemmas 26, 27, 28, 29, 30, 31, 32 and 33 guarantee that there exists a desired set of edges $F$ of constant size.
To bound the total running time including recursive steps, let $T(\sigma)$ be the running time of Algorithm 1 as a function of $\sigma$. In each step, starting with a problem of size $\sigma$, we generate either a single subproblem of size $\sigma^{\prime}<\sigma$ (to easily see this in Algorithm 2 recall that $7\leq|V_{1}|\leq|V_{2}|$) or two subproblems of size $\sigma^{\prime}$ and $\sigma^{\prime\prime}$ such that $\sigma^{\prime},\sigma^{\prime\prime}<\sigma$ and $\sigma^{\prime}+\sigma^{\prime\prime}\leq\sigma$. Thus, $T(\sigma)\leq\max\{T(\sigma^{\prime}),T(\sigma^{\prime})+T(\sigma^{\prime\prime})\}+\mathrm{poly}(\sigma)$ with $\sigma^{\prime},\sigma^{\prime\prime}<\sigma$ and $\sigma^{\prime}+\sigma^{\prime\prime}\leq\sigma$. Therefore, $T(\sigma)\leq\sigma\cdot\mathrm{poly}(\sigma)$. ∎
Lemma 35 (correctness).
Proof.
The proof is by induction over the tuple $(\lvert V(G)\rvert,\lvert E(G)\rvert)$. The base cases are given by Algorithms 1 and 1 of Algorithm 1. If we enter Algorithm 1, the claim trivially holds. If we call $\mathsf{ALG}$ in Algorithm 1, the claim follows by the correctness of $\mathsf{ALG}$ for $(\alpha,\varepsilon)$-structured instances. If in Algorithm 1 any of the conditions Algorithms 1, 1, 1, 1 and 1 apply, the claim follows from [18, Lemma A.6].
Thus, for the remaining proof we consider the case where Algorithm 1 holds and we enter Algorithm 2. In this case the correctness follows from the induction hypothesis, Corollary 23 and Lemmas 27, 28, 29, 30, 31, 32 and 33. ∎
Let $\mathsf{red}(G)\coloneq|\mathsf{RED}(G)|$ denote the size of the solution computes by $\mathsf{RED}$ for a given 2-ECSS instance $G$. The following lemma implies $\mathsf{red}(G)\leq(\alpha+4\varepsilon)\mathsf{opt}(G)$ using the trivial bound $\mathsf{opt}(G)\geq|V(G)|$. While its proof is quite long, it is not difficult and easy to verify: We consider different cases depending on the reduction the algorithm uses and, if Algorithm 2 is executed, we distinguish for some cases further which solution types the actual optimal solution uses for both sides of the cut. For each case, we can easily derive bounds for the different parts of the solution that our algorithm produces, and put them together to obtain the overall bound, which is stated below.
Lemma 36 (approximation guarantee).
For any 2-ECSS instance $G$ it holds that
Proof.
The proof is by induction over the tuple $(\lvert V(G)\rvert,\lvert E(G)\rvert)$. The base cases are given by Algorithms 1 and 1 of Algorithm 1. If we enter Algorithm 1, the stated bound trivially holds. If we call $\mathsf{ALG}$ in Algorithm 1, the claim follows because we assume that $\mathsf{ALG}(G)$ returns an $\alpha$-approximate solution to the 2-ECSS problem on $G$ and $G$ is in this case $(\alpha,\varepsilon)$-structured by the correctness of $\mathsf{RED}$ (cf. Lemma 35).
If in Algorithm 1 any of the reductions in Algorithms 1, 1, 1, 1 and 1 is executed, the claim follows from [18, Lemma A.7]. (The proof in [18] proves the reduction with even stronger assumptions, which can be relaxed to our constants.)
If Algorithm 2 is executed, let $\mathsf{OPT}_{i}\coloneq\mathsf{OPT}(G)\cap E(G_{i})$ and $\mathsf{opt}_{i}\coloneq\lvert\mathsf{OPT}_{i}\rvert$ for $i\in\{1,2\}$. Note that $\mathsf{opt}(G)=\mathsf{opt}_{1}+\mathsf{opt}_{2}$. For the remainder of this proof, we only consider the execution of Algorithm 2. In particular, all line references (if not specified differently) refer to Algorithm 2. First, suppose that the condition in Algorithm 2 holds. Note that for $i\in\{1,2\}$ we have that $\mathsf{OPT}_{i}$ is a feasible solution for $G_{i}^{\prime}$ and therefore $\mathsf{opt}_{i}\geq\mathsf{opt}(G_{i}^{\prime})$. Using the induction hypothesis, Lemma 26 and $\lvert V(G_{1})\rvert+\lvert V(G_{2})\rvert=\lvert V_{1}\rvert+\lvert V_{2}\rvert+2=\lvert V(G)\rvert-1$, we conclude that
@@
It remains to verify the case where we reach Algorithm 2. We distinguish several cases depending on which case the algorithm uses. Note that Lemma 21 gives $\mathsf{opt}_{1}\geq\mathsf{opt}_{1}^{\min}\geq 8$. Furthermore, throughout the proof we require $\alpha\geq\frac{5}{4}$. We use both statements in every case below.
@@
Note that $\mathsf{opt}(G_{2}^{\mathsf{B1}})\leq\mathsf{opt}_{2}$, because $u$, $v$, and $w$ are contracted into a single vertex in $G_{2}^{\mathsf{B1}}$. Thus, $\mathsf{opt}(G)=\mathsf{opt}_{2}+\mathsf{opt}_{1}\geq\mathsf{opt}(G_{2}^{\mathsf{B1}})+\mathsf{opt}_{1}$. Moreover, $\mathsf{opt}_{1}^{\mathsf{B1}}\leq\mathsf{opt}_{1}^{\min}+1\leq\mathsf{opt}_{1}+1$, and $\lvert F^{\mathsf{B1}}\rvert\leq 1$ by Lemma 27. Therefore, we conclude
@@
(Algorithms 2 and 2: $t^{\min}=\mathsf{B2}$ or $t^{\min}=\mathsf{C1}$)
Let $\tau\coloneq t^{\min}$. Note that $\mathsf{opt}(G_{2}^{\tau})\leq\mathsf{opt}_{2}$, because $u$, $v$, and $w$ are contracted into a single vertex in $G_{2}^{\tau}$. Thus, $\mathsf{opt}(G)\geq\mathsf{opt}_{2}+\mathsf{opt}_{1}(G_{1}^{\tau})\geq\mathsf{opt}(G_{2}^{\tau})+\mathsf{opt}_{1}(G_{1}^{\tau})$. Moreover, $\lvert F^{\tau}\rvert\leq 2$ by Lemmas 28 and 29. Therefore, we conclude
@@
In Algorithm 2 we have the following $\mathsf{C2}$ subtypes, which use different reductions $G_{2}^{\mathsf{C2}}$ for the recursion (Algorithms 2, 2 and 2).
@@
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains a $C(u)-C(v)$ path. In this case the definition of Algorithm 2, Lemma 30, and the induction hypothesis gives
@@
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains either a $C(u)-C(v)$ path or a $C(v)-C(w)$ path. In this case the definition of Algorithm 2, Lemma 31, and the induction hypothesis gives
@@
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains some path between $C(u)$, $C(v)$, and $C(w)$. In this case the definition of Algorithm 2, Lemma 32, and the induction hypothesis gives
@@
We distinguish the following cases depending on which type combination of Proposition 18 is present for $(\mathsf{OPT}_{1},\mathsf{OPT}_{2})$.
@@
    $(\mathsf{A},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, Lemma 22 gives $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\mathsf{C2}}+3$.
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,vw\}$ and at most two more edges, which are guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (1) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+6$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy,wy\}$ and at most two more edges, which are guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,wy\}$ and at most two more edges, which are guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives
@@
    $(\mathsf{B1},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{B1}}$ exists but the condition in Algorithm 2 did not apply, we conclude that $\mathsf{opt}_{1}^{\mathsf{B1}}\geq\mathsf{opt}_{1}^{\min}+2=\mathsf{opt}_{1}^{\mathsf{C2}}+2$.
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,vw\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (1) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy,wy\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,wy\}$ and at most one more edge, which is guaranteed by Lemma 24 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives

### v2

@@ A.2.2 Auxiliary Lemmas for Algorithm˜3
If $C_{2}^{\mathsf{C3}}(v)$ is connected to $C_{2}^{\mathsf{C3}}(w)$ in $H_{2}^{\mathsf{C3}}$, then $H_{2}^{\mathsf{C3}}$ is connected. Further, $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is connected and spanning for $G_{1}$. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). We conclude via Lemma˜36 that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-3\leq 0$.
Otherwise, that is, $C_{2}^{\mathsf{C3}}(v)$ is not connected to $C_{2}^{\mathsf{C3}}(w)$ in $H_{2}^{\mathsf{C3}}$, Lemma˜35 guarantees the existence of an edge $e$ in $G_{2}$ between the vertices of $V(C_{2}^{\mathsf{C3}}(w))$ and the vertices of $V(H_{2}^{\mathsf{C3}})\setminus V(C_{2}^{\mathsf{C3}}(w))$. Therefore, $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Hence, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=3-3=0$.
@@
First, assume that $H_{2}^{\mathsf{C3}}$ is connected. We have that $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is connected and spanning for $G_{1}$. Thus, every edge in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ (resp. $H_{2}^{\mathsf{C3}}$) is part of a 2EC component of $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) or lies on a path in $\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ ($H_{2}^{\mathsf{C3}}$) between any two distinct vertices of $\{u,v,w\}$ between which is a path in $H_{2}^{\mathsf{C3}}$ ($\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$). We conclude via Lemma˜36 that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=2-4\leq 0$.
@@
If $H_{2}^{\mathsf{C3}}$ is composed of two connected components, assume w.l.o.g. that $C_{2}^{\mathsf{C3}}(u)$ is connected to $C_{2}^{\mathsf{C3}}(v)$ in $H_{2}^{\mathsf{C3}}$ and the other connected component in $H_{2}^{\mathsf{C3}}$ is $C_{2}^{\mathsf{C3}}(w)$. In this case, Lemma˜35 guarantees that there is an edge $e$ in $E(G_{2})\setminus C_{2}^{\mathsf{C3}}(w)$ between the vertices of $C_{2}^{\mathsf{C3}}(w)$ and the vertices of $V(H_{2}^{\mathsf{C3}})\setminus V(C_{2}^{\mathsf{C3}}(w))$. Therefore, $H_{2}^{\mathsf{C3}}\cup\{e\}$ is connected. Hence, we can analogously to the previous case show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq\{e,e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$, and $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=3-4\leq 0$.
@@
Otherwise, that is, $C_{2}^{\mathsf{C3}}(u)$, $C_{2}^{\mathsf{C3}}(v)$ and $C_{2}^{\mathsf{C3}}(w)$ are all pairwise disjoint, Lemma˜35 guarantees the existence of two edges $F\subseteq E(G_{2})$ such that $H_{2}^{\mathsf{C3}}\cup F$ is connected. Hence, we can analogously to the previous case(s) show that $H_{2}^{\mathsf{C3}}\cup\mathsf{OPT}_{1}^{\mathsf{C3}}\cup F^{\mathsf{C3}}$ with $F^{\mathsf{C3}}\coloneq F\cup\{e^{\mathsf{C3}}_{uv},e^{\mathsf{C3}}_{vw}\}$ is a 2EC spanning subgraph of $G$. We have that $|F^{\mathsf{C3}}|-|\mathsf{RED}(G_{2}^{\mathsf{C3}})|+|H_{2}^{\mathsf{C3}}|=4-4=0$. ∎
Finally, we prove that the reductions of Algorithm˜3 preserve the approximation factor. While the proof is quite long, it is of the same spirit of the analogue for Algorithm˜2 (cf. Lemma˜30) and easy to verify: We consider different cases depending on the reduction the algorithm uses and, if Algorithm˜3 is executed, we distinguish for some cases further which solution types the actual optimal solution uses for both sides of the cut. For each case, we can easily derive bounds for the different parts of the solution that our algorithm produces, and put them together to obtain the overall bound, which is stated below.
Lemma 45.
Let $G$ be a 2-ECSS instance. If every recursive call to $\mathsf{RED}(G^{\prime})$ in Algorithm˜3 on input $G$ satisfies $\mathsf{red}(G^{\prime})\leq\alpha\cdot\mathsf{opt}(G^{\prime})+4\varepsilon\cdot|V(G^{\prime})|-4$, then it holds that $\mathsf{red}(G)\leq\alpha\cdot\mathsf{opt}(G)+4\varepsilon\cdot|V(G)|-4$.
@@
Let $\mathsf{OPT}_{i}\coloneq\mathsf{OPT}(G)\cap E(G_{i})$ and $\mathsf{opt}_{i}\coloneq\lvert\mathsf{OPT}_{i}\rvert$ for $i\in\{1,2\}$. Note that $\mathsf{opt}(G)=\mathsf{opt}_{1}+\mathsf{opt}_{2}$. For the remainder of this proof, all line references (if not specified differently) refer to Algorithm˜3. Note that Lemma˜33 gives $\mathsf{opt}_{1}\geq\mathsf{opt}_{1}^{\min}\geq 8$. Furthermore, throughout the proof we require $\alpha\geq\frac{5}{4}$. We distinguish in the following which reduction the algorithm uses.
(Algorithm 3: $|V_{1}|>\frac{2}{\varepsilon}-4$)
Note that for $i\in\{1,2\}$, we have that $\mathsf{OPT}_{i}$ is a feasible solution for $G_{i}^{\prime}$ and therefore $\mathsf{opt}_{i}\geq\mathsf{opt}(G_{i}^{\prime})$. Using the induction hypothesis, Lemma˜37 and $\lvert V(G_{1})\rvert+\lvert V(G_{2})\rvert=\lvert V_{1}\rvert+\lvert V_{2}\rvert+2=\lvert V(G)\rvert-1$, we conclude that
@@
Note that $\mathsf{opt}(G_{2}^{\mathsf{B1}})\leq\mathsf{opt}_{2}$, because $u$, $v$, and $w$ are contracted into a single vertex in $G_{2}^{\mathsf{B1}}$. Thus, $\mathsf{opt}(G)=\mathsf{opt}_{2}+\mathsf{opt}_{1}\geq\mathsf{opt}(G_{2}^{\mathsf{B1}})+\mathsf{opt}_{1}$. Moreover, $\mathsf{opt}_{1}^{\mathsf{B1}}\leq\mathsf{opt}_{1}^{\min}+1\leq\mathsf{opt}_{1}+1$, and $\lvert F^{\mathsf{B1}}\rvert\leq 1$ by Lemma˜38. Therefore, we conclude
@@
(Algorithms 3 and 3: $t^{\min}=\mathsf{B2}$ or $t^{\min}=\mathsf{C1}$)
Let $\tau\coloneq t^{\min}$. Note that $\mathsf{opt}(G_{2}^{\tau})\leq\mathsf{opt}_{2}$, because $u$, $v$, and $w$ are contracted into a single vertex in $G_{2}^{\tau}$. Thus, $\mathsf{opt}(G)\geq\mathsf{opt}_{2}+\mathsf{opt}_{1}(G_{1}^{\tau})\geq\mathsf{opt}(G_{2}^{\tau})+\mathsf{opt}_{1}(G_{1}^{\tau})$. Moreover, $\lvert F^{\tau}\rvert\leq 2$ by Lemmas˜39 and 40. Therefore, we conclude
@@
In Algorithm˜3 we have the following $\mathsf{C2}$ subtypes, which use different reductions $G_{2}^{\mathsf{C2}}$ for the recursion (Algorithms˜3, 3 and 3).
@@
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains a C(u)C(v)C(u)-C(v) path. In this case the definition of Algorithm˜3, Lemma˜41, and the induction hypothesis gives
@@
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains either a C(u)C(v)C(u)-C(v) path or a C(v)C(w)C(v)-C(w) path. In this case the definition of Algorithm˜3, Lemma˜42, and the induction hypothesis gives
@@
Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains some path between $C(u)$, $C(v)$, and $C(w)$. In this case the definition of Algorithm˜3, Lemma˜43, and the induction hypothesis gives
@@
We distinguish the following cases depending on which type combination of Proposition˜17 is present for $(\mathsf{OPT}_{1},\mathsf{OPT}_{2})$.
@@
    $(\mathsf{A},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, Corollary˜34 gives $\mathsf{opt}_{1}=\mathsf{opt}_{1}^{\mathsf{A}}\geq\mathsf{opt}_{1}^{\mathsf{C2}}+3$.
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,vw\}$ and at most two more edges, which are guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (1) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+6$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy,wy\}$ and at most two more edges, which are guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,wy\}$ and at most two more edges, which are guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives
@@
    $(\mathsf{B1},\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2}\})$. Since $\mathsf{OPT}_{1}^{\mathsf{B1}}$ exists but the condition in Algorithm˜3 did not apply, we conclude that $\mathsf{opt}_{1}^{\mathsf{B1}}\geq\mathsf{opt}_{1}^{\min}+2=\mathsf{opt}_{1}^{\mathsf{C2}}+2$.
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,vw\}$ and at most one more edge, which is guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, continuing (1) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+5$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vz,zy,wy\}$ and at most one more edge, which is guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (2) gives
@@
    We have $\mathsf{opt}(G_{2}^{\mathsf{C2}})\leq\mathsf{opt}_{2}+4$ because we can turn $\mathsf{OPT}_{2}$ with the dummy edges $\{uy,vy,wy\}$ and at most one more edge, which is guaranteed by Lemma˜35 if required, into a solution for $G_{2}^{\mathsf{C2}}$. Thus, (3) gives